OpenStudy (anonymous):

a certain tennis player makes a successful first serve 71% of the time. assume that each serve is independent of the others. if she serves 5 times, what's the probability she gets c) at least 4 serves in? d) no more than 4 serves in?

OpenStudy (anonymous):

tk help me with one after this?

OpenStudy (tkhunny):

Think about the probability that NO serve makes it in. p(miss)*p(miss)*p(miss)*p(miss)*p(miss) = 0.29^5 What does that do for us?

OpenStudy (anonymous):

what? I dont understand

OpenStudy (tkhunny):

Ignoring the problem statement for a moment, what is the probability that ALL FIVE serves will be clinkers?

OpenStudy (anonymous):

.180

OpenStudy (anonymous):

I understand how to do the first two parts. I need help setting up the formula for c and d please

OpenStudy (anonymous):

Al teast 4 means 4 or 5 in this case.

OpenStudy (tkhunny):

No, that is the probability that all five will be good. The probability that all 5 will be bad is 0.29^5 = 0.00205 Now, we have: p(0 good) = 0.00205 p(1 good) = p(2 good) = p(3 good) = p(4 good) = p(5 good) = 0.18042 That's just calculating the easy ones. Does this enable us to answer the questions? c) at least 4 serves in? p(at least 4) = p(4) + p(5) -- Okay, so we don't quite know enough for this one. d) no more than 4 serves in? p(no more than 4) = p(0) + p(1) + p(2) + p(3) + p(4) Aha!!! We CAN solve this one. p(no more than 4) = 1 - p(5) -- Do you see this?

OpenStudy (anonymous):

No more than 4 would mean 0, 1, 2, 3, or 4

OpenStudy (anonymous):

okay, got it thanks